What is the NPV of a project which reauires an initial investment od $100,000, generates cash flows kf $25,000, $50,000, $75,000, and $100,000 in years 1 through 4, then requires another investment of $50,000 in year 5? The required rate of return is 12%.

A. $100,000
B. $29,116
C. $50,745
D. $79,116
Answer #1

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=25000/1.12+50,000/1.12^2+75000/1.12^3+100,000/1.12^4

=$179116.45

Present value of outflows=100,000+50,000/1.12^5

=128371.34

NPV=Present value of inflows-Present value of outflows

=179116.45-128371.34

=$50745(Approx).